Given the demand for a product as Qd=100-5P and the supply is given as Qs=-20+3P. You are told equilibrium is obtained at the point where Qd=Qs. The equilibrium price for the product is

A. R5

B. R6.67

C. R15

D. R20


1
Expert's answer
2021-06-21T11:57:29-0400
"Q_d=Q_s,""100-5P_E=-20+3P_E,""120=8P_E,""P_E=R15."

Therefore, the correct answer is (C) R15.